Fermi National Laboratory


Accelerator Update for January 31, 2005 - February 2, 2005

Monday January 31
The day shift began with the Tevatron (TeV) in shot setup and with MiniBooNE taking beam.

Operations established store 3953 at 8:23 AM with an initial luminosity of 85.46E30. The Antiproton Source (Pbar) resumed stacking.

Operations reported at 8:36 AM that the Switchyard (SY) vacuum was fixed.

At 10:35 AM, Operations reported loosing the Tevatron Electron Lens (TEL) and a Pbar RF station (ARF1). Fortunately, the stack and the store survived. The High Voltage expert said that ComEd reported a power glitch.

Fixed target experts began tuning beam to the Meson experiments MTest and MCenter at 11:30 AM.

Operations put the Recycler into shot setup and began stashing antiprotons at 1:32 PM.

Operations completed a Pbar reverse proton tune-up at 2:17 PM. Linac and Booster technicians began a number of maintenance jobs. This halted stacking and SY beam.

Pbar resumed stacking at 3 PM.

At 4:22 PM, power supply experts began replacing a Main Injector (MI) sextupole power supply. This halted stacking and SY beam.

Pbar resumed stacking 5:05 PM.

Tuesday February 1
The midnight shift began with Operations monitoring store 3953 and the stash, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

At 12:56 AM, a TeV RF station (TRF7) tripped off due to its water heater tripping off. By 1:45 AM, this was causing a serious impact on luminosity and lifetimes.

At 1:56 AM, the MI safety system beam permit dropped due to a power supply (I:Q703) trip. Operators couldn't reset the supply. At 2:29 AM, operators called in expert help. However, before the expert arrived operators had managed to reset the supply at 2:53 AM.

At 5:04 AM, the MiniBooNE horn tripped off. It tripped again at 6:45 AM. The horn reset okay each time.

Operations terminated store 3953 at 12:12 PM and then put the TeV into shot setup about fifteen minutes later.

At 1 PM, Operations discovered the MI coalescing cavities on and at a 20 kV level, which had heated them up. MI experts turned them off to cool. This delayed shot setup.

At 1:50 PM, cryo system techs at CHL reported that the TeV sector F4 wet engine speed wouldn't regulate. Experts investigated. This also delayed the shot setup.

Cryo system techs reported at 3:45 PM that they had found and replaced a bad digital-to-analog converter card for the F4 wet engine.

The Recycler stashed antiprotons around 4 PM. Operators also put the TeV through a dry squeeze around 4 PM.

At 4:57 PM, Operations prepared the TeV and MI for an access into the TeV F sector to check the F17 kicker. The Collider Detector at Fermilab (CDF) and D-Zero (D0) accessed their collision halls during this F17 access, which occurred around 6 PM.

Operations reestablished beam to MiniBooNE at 6:49 PM.

Operations began the TeV shot setup 9:18 PM.

Operations established store 3956 at 11PM with an initial luminosity of 94.9E30. Pbar resumed stacking.

Operations began sending beam to MCenter at 11:22 PM.

Wednesday February 2
The midnight shift began with Operations monitoring store 3956, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the Recycler into shot setup at 3:31 AM. This halted all SY beam.

The Recycler finished stashing at 5:30 AM. Pbar resumed stacking.

MCenter resumed taking beam at 5:51 AM.

The Plan for Wednesday
The plan for today is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/2/2005   email Fermilab